OpenStudy (anonymous):

What is the rule when multipling and dividing with integers? 6/-1 or 6 x -7

OpenStudy (bahrom7893):

what do u mean? 6/-1 = -6; 6 * -7 = -42

OpenStudy (anonymous):

solve 6x-7 or it could be a division problem,\. What are the rules to follow

OpenStudy (bahrom7893):

Solve: 6x - 7 = SOMETHING?

OpenStudy (bahrom7893):

Is x a letter or multiplication sign?

OpenStudy (anonymous):

multiplication sign

OpenStudy (bahrom7893):

There are no rules to follow.. you just multiply and divide, BUT: if u are multiplying or dividing - number by + number, ur result will be a - number if u multiply or divide a + by + or - by -, you will get plus... Same signs give u plus, diff signs, minus

OpenStudy (anonymous):

-7 x 7= -49 35 divided by -7=-5 -32 divided by -8=4 so are these correct

OpenStudy (bahrom7893):

YUP U GOT EM!!!

OpenStudy (anonymous):

thanks thats all I need

OpenStudy (bahrom7893):

Click on become a fan, please? I wanna reach 60, thanks!